# Outcome-based Reinforcement Learning to Predict the Future

### v3

@@ 4 Conclusion
This paper demonstrates the effective application of Reinforcement Learning with Verifiable Rewards (RLVR) to the challenging domain of forecasting real-world events. We train a variety of reasoning models that take as as input a TRUE/FALSE forecasting question, alongside a set of news articles published before the prediction date, and aims to predict how the question will resolve. We compare a variety of training algorithms for reasoning models, using a sample of 100k historical questions (10k real + 90k synthetic).
On a test set of recent forecasting questions, our best-performing model was a 7-run ensemble trained using ReMax, and with guardrails to improve learning stability. This model reached a brier score of 0.190: significantly outperforming much larger frontier models. We find that these gains are also practically meaningful: in a trading simulation, the ReMax ensemble earned a total of \$52 in profit from simulated bets on Polymarket, with a 10% return on investment - or 20% among questions with low market confidence. These results highlight the promise of well-calibrated RL-trained forecasters to serve as tools to support human reasoning under uncertainty.

### v4

@@ 4 Conclusion
This paper demonstrates the effective application of Reinforcement Learning with Verifiable Rewards (RLVR) to the challenging domain of forecasting real-world events. We train a variety of reasoning models that take as input a TRUE/FALSE forecasting question, alongside a set of news articles published before the prediction date, and aims to predict how the question will resolve. We compare a variety of training algorithms for reasoning models, using a sample of 110k historical questions (10k from Polymarket + 100k generated).
On a test set of recent forecasting questions, our best-performing model was a 7‑run ensemble trained using ReMax, and with guardrails to improve learning stability. This model reached a Brier score of 0.190: outperforming much larger frontier models. We find that these gains are also practically meaningful: in a trading simulation, the ReMax ensemble earned a total of \$52 in profit from simulated bets on Polymarket, with a 10% return on investment - or 20% among questions with low market confidence. To be clear, we do not intend for this simulation to be understood as an estimate for the impact our model may have on real-world financial trading: Financial markets operate under different constraints to prediction markets and, among other differences, are many thousands of times larger in volume than prediction markets. Rather, our results highlight the promise of well-calibrated RL-trained forecasters to serve as tools to support human reasoning under uncertainty.
5 Broader Impact
Our results carry potential societal implications and risks. For one, while we do not directly evaluate our approach in the domain of financial trading, it is realistic that our model or training strategy may be employed by individuals or institutions to guide investment in financial markets. Furthermore, prediction markets themselves frequently inform societal decisions in many high stakes domains such as policy. In both cases, if an AI model were used directly to make high stakes decisions, then inaccuracies or systematic biases in the model could carry substantial risks. However, we consider it a strength of our method that it provides human-readable chain-of-thought traces to accompany its predictions, which greatly aid interpretability. In general, we expect that the best use of such models is as tools to support, rather than replace, existing processes of human decision-making.
[…]